# Share the Sugar

### v1

@@ 7 Responses
So there is reason to consider giving up either Personal Good or Negative Dominance. We won’t have much new to say about the former possibility, but will mention a few familiar considerations that might push one in that direction. [^24] In conversation, Ralf Bader (who independently noticed the central puzzle of this paper) says that he would give up Personal Good. First, as noted earlier, egalitarians may reject Personal Good for reasons unrelated to incompleteness, since it requires insensitivity to distributive features of outcomes that egalitarians think are important. For example, in the following table, Personal Good prefers the lottery on the right, while ex post egalitarians have reason to prefer the lottery on the left, since it guarantees perfect equality. [^25] It’s not obvious how much this point undercuts the preceding arguments, though, since we can plausibly adapt those arguments to guarantee that the outcome preferred by Personal Good involves equal or greater ex post equality. In Table 2, for instance, this will be the case if we sour an outcome in Lottery 1 rather than sweetening an outcome in Lottery 2. Guaranteeing that we can do this, though, requires a modest strengthening of Small Improvements.
Heads Tails
$P_{1}$ $1$ 00
$P_{2}$ $1$ 00
Heads Tails
$P_{1}$ $1+\epsilon$ 00
$P_{2}$ 00 $1+\epsilon$
Second, as already noted, you might be willing to reject Personal Good because of its incompatibility with dominance principles in infinite contexts. Finally, while the idea that “what is at least as good for everyone, and better for some, is better overall” is extremely compelling, it is not obvious whether a lottery can be “good for” an individual—we derive no benefit from how our lives might have turned out, over and above how they in fact turn out. The “ex post individualism” expressed by Pareto is, in this respect, more compelling than the “ex ante individualism” expressed by Personal Good.
None of these considerations, though, have any obvious connection with incompleteness. So the desire to preserve incompleteness in the face of the results from the last two sections does not seem to provide any distinctive motivation for jettisoning Personal Good. It would be interesting if it turns out to be possible to find such motivations. Perhaps, for instance, there is a view which endorses population-level incompleteness without endorsing incompleteness for individual lives, and which sees incompleteness as arising from global distributive features of outcomes whose significance counts against Personal Good in something like the way that egalitarian concerns do. But we won’t try to develop such a view here.
A route that is more directly motivated by incompleteness, we think, is to reject Negative Dominance. [^26] One view that takes this route is Gustafsson’s (2020) Undistinguished Critical-Range Utilitarianism (see note 17). Although he does not discuss this point explicitly, Gustafsson’s preferred rule for evaluating risky prospects (p. 94) satisfies Personal Good but not Negative Dominance. Someone who accepts Personal Good might motivate the rejection of Negative Dominance on “individualist” grounds as follows: “As individualists, our fundamental concern is with the personal good of each individual considered separately. ‘Impersonal value’ is just a sort of summary statistic. For instance, the claim that $(2,0)$ is impersonally better than $(0,1)$ is a useful but incomplete way of summarizing more fundamental facts, namely, that $(2,0)$ is better for $P_{1}$, worse for $P_{2}$, and the first difference in personal value is greater than the second. The concept of impersonal value plays a useful role—for instance, it is an appropriate guide to action under certainty. But because it doesn’t capture everything we fundamentally care about, comparisons of lotteries based only on the impersonal value of their possible outcomes—like the comparisons expressed by dominance principles—will sometimes overlook important considerations. These oversights show up in infinite contexts, where dominance reasoning can recommend options that are worse for all concerned. What the arguments of §§45 show is that similar oversights can happen in finite contexts: Negative Dominance, because it is sensitive only to the ‘overall’ value of outcomes, can miss a decisive reason for ranking one lottery above another.” This strikes us as at least a plausible reason why sufficiently committed individualists might not feel attached to Negative Dominance.
A second point is that we can soften the blow of rejecting Negative Dominance, by replacing it with a weaker principle that captures some of its intuitive appeal:
@@
Lottery $L_{1}$ is better than lottery $L_{2}$ only if, on every admissible completion of personal and impersonal betterness, some possible outcome $o_{1}$ of $L_{1}$ is better than some possible outcome $o_{2}$ of $L_{2}$.
An “admissible completion” is one that satisfies some set of fundamental constraints on the ranking of outcomes. Suppose that these constraints include transitivity and (ex post) Pareto. Then we conjecture that Supervaluational Negative Dominance will be compatible, in finite contexts, with the various sets of assumptions that were shown to conflict with Negative Dominance in §§45.
For instance, in the case in Table 2, every admissible completion of betterness rankings of lives and outcomes must rank either $(a^{+},a)$ or $(b,b)$ above $(a,b)$ and $(b,a)$. [^27] Suppose that, on a given completion, $a$ is at least as good as $b$ in terms of personal value. Then, by transitivity, $a^{+}$ is better than $b$, and by Pareto, $(a^{+},a)$ is strictly better than $(a,b)$ and $(b,a)$. The other possibility is that $b$ is strictly better than $a$, in which case $(b,b)$ is strictly better than $(a,b)$ and $(b,a)$ by Pareto.
Arguably, the intuitive appeal of Negative Dominance is explained by its similarity to the slightly subtler principle of Supervaluational Negative Dominance. This conclusion will be especially appealing to those who understand incompleteness in terms of indeterminacy between complete orderings (e.g. Broome (1997), Dorr et al. (2021)). On this understanding, Negative Dominance says that for $L_{1}$ to be better than $L_{2}$, there must be some outcome of $L_{1}$ that is determinately better than some outcome of $L_{2}$; but all that is really required is that it’s determinately the case that some outcome of $L_{1}$ is better than some outcome of $L_{2}$—it can be indeterminate which pair of outcomes those are.
It is worth noting, however, that Supervaluational Negative Dominance still conflicts with Personal Good in infinite contexts (e.g., in Table 7 above). So the retreat to this weaker principle holds less appeal for those whose rejection of Negative Dominance is motivated by an absolute commitment to individualistic principles like Personal Good.
Whatever its motivation, the rejection of Negative Dominance would be a surprising and significant conclusion. It’s an intuitively compelling principle in its own right. But also, as we have seen, many fans of incompleteness (like Schoenfield (2014), Bales et al. (2014), and Doody (2023)) have adopted the stronger Statewise Negative Dominance as a central principle for evaluating risky options with incomparable outcomes. If the preceding arguments force us to give up Negative Dominance then, a fortiori, they force us to reject this family of views as well.
@@ 8 Conclusion
Against weak background assumptions, Personal Good and Negative Dominance rule out many intuitive forms of value incomparability. Determining which of these should be rejected will require making progress on deep and pressing questions. To mention a few salient ones: is the betterness of lives and outcomes conceptually prior to the betterness of lotteries, in such a way that we cannot reason from principles about lotteries to conclusions about the comparative betterness of lives or outcomes? Is there a distinctive “impersonal” notion of betterness – betterness from the perspective of the world – that can’t be reduced to facts about how things go for particular individuals? How should we understand phenomena which putatively support the existence of value incomparability?
Our incompatibility results display surprising constraints on the answers to these questions, and connections between them. But far from being resisted, we think these constraints should be welcomed. They illuminate the difficult choices that must be made in developing a general theory of the good, and thus may point the way toward more satisfactory theories.

### v2

@@ 7 Objections
Table 9 illustrates this inconsistency. The table describes a pair of lotteries in which, for every possible individual, the personal lottery given by Lottery 1 stochastically dominates the personal lottery given by Lottery 2; and yet Lottery 2 is certain to result in a positive total of personal value, while Lottery 1 is certain to result in a negative total of personal value. [^28] This example builds on one in Kowalczyk (forthcoming). Related examples of conflicts between dominance principles and ex ante Pareto-like principles in infinitary contexts appear in Goodsell (2021), Wilkinson (2023), Nebel (2023), and Hong and Russell (forthcoming). Personal Good implies that Lottery 1 is better than Lottery 2. But it is plausible that every outcome of Lottery 2 is better than every outcome of Lottery 1, in which case Negative Dominance implies that Lottery 1 cannot be better. This latter claim would follow straightforwardly given a total utilitarian theory of impersonal value, but it follows from many other views as well and is in any case intuitively plausible.292929 Totalism provides a simple illustration of the argument in the main text, but the axiological assumptions required to generate this sort of conflict between Negative Dominance and Personal Good are very weak. It would suffice, for instance, to assume a principle we might call “Inverse Lives”: There exists a pair of lives $l^{+}$ and $l^{-}$ such that, among outcomes in which everyone who exists has one of those two lives, any outcome where most people have $l^{+}$ is better than any outcome where most people have $l^{-}$. This assumption is endorsed, for instance, by standard versions of total utilitarianism, critical-level utilitarianism, prioritarianism, average utilitarianism, variable-value utilitarianism, and many forms of egalitarianism. We have presented the example as involving only finite (though unboundedly large) populations. We could switch to an infinite-population example by replacing every instance of nonexistence in Table 9 with a neutral life, with welfare 00. This would allow us to weaken Personal Good by restricting it to fixed-population contexts (where the same individuals exist in every outcome of both lotteries). Then the principle could simply assert that if one lottery gives every individual a stochastically dominant personal lottery, then it’s impersonally better. While it is not obvious which principle should be given up in this case, it shows that Negative Dominance and Personal Good cannot both be unrestrictedly true. And if a prima facie plausible principle must be abandoned in one context, that significantly undercuts its appeal in other contexts as well.
Response
We reject the premise that needing to give up a principle in infinite contexts must significantly undercut its appeal in finite contexts. [^30] For discussion of this issue, see for instance Russell (2023) and Kowalczyk (forthcoming). There are many principles that hold in finite but not infinite domains—for instance, famously, the principle that no set can be put into one-to-one correspondence with any of its proper subsets. As a closer analogy, although Pareto and the standard Anonymity principle (which says that permuting a welfare distribution in an outcome results in an equally good outcome) are incompatible for infinite populations (Van Liedekerke 1995), most ethicists and social welfare theorists accept both of these principles for finite populations, and use them happily as premises in that context. So although either Negative Dominance or Personal Good must be given up in certain infinite contexts, this need not seriously weaken the case for completeness in finite ones. In particular, Negative Dominance and Personal Good are perfectly consistent when restricted to lotteries with finitely many possible outcomes, even when some of those outcomes involve infinite populations. [^31] Here’s a theory of value that satisfies both principles for infinite-population, finite-support lotteries: Let the set of possible individuals be indexed by the natural numbers, and assume a free ultrafilter $U$ on the naturals. Suppose further that the personal betterness ordering of lives is complete, and that the value of lives (including nonexistence) can be represented by real numbers. Let $\operatorname{EV}(P_{i}|L_{j})$ be the expected personal value for individual $P_{i}$ under lottery $L_{j}$. Then the following betterness relation on lotteries will satisfy both Negative Dominance and Personal Good: $L_{1}$ is at least as impersonally good as $L_{2}$ if and only if the set of $n$ such that $\sum_{i=0}^{n}\operatorname{EV}(P_{i}|L_{1})\geq\sum_{i=0}^{n}\operatorname{EV}(P_{i}|L_{2})$ belongs to the ultrafilter $U$. This view satisfies Personal Good unrestrictedly and satisfies Negative Dominance for finite-support lotteries. An implication of this consistency is that our result in §5 ruling out strong incomparability between distributions could be extended to infinite distributions—ruling out incomparability between, for instance, the distributions $(1,0,1,0,\ldots)$ and $(0,1,0,1,\ldots)$. But we have some reservations about the application of Personal Good to infinite-population, finite-support lotteries, as we explain in the next note. The proofs of our results in §§46 used only lotteries with finite support, and so don’t require Negative Dominance or Personal Good to hold outside that context.323232 Then why restrict ourselves to finite populations, as well as finitely supported lotteries? With respect to our result in §6 (concerning the ranking of lotteries), the main reason is that the principle of Substitution is dubious in infinite-population contexts (see n. 25). With respect to the result in §5 (concerning the ranking of outcomes), the story is slightly more subtle. First, we currently incline to the view that the principle to give up in Table 9 is Personal Good. Notably, in this example, Personal Good conflicts with just about any imaginable dominance principle, including Stochastic Dominance, Statewise Dominance, and even Superdominance—the principle that, if each possible outcome of $L_{1}$ is better than every possible outcome of $L_{2}$, then $L_{1}$ is better than $L_{2}$. Restricting Negative Dominance therefore wouldn’t do us any good here unless we were also willing to restrict those other dominance principles. And those principles are, intuitively, absolutely essential to the idea that the final value of outcomes should be our guide to assessing lotteries. On the other hand, we can tell a plausible story about why Personal Good should fail here: A stochastic improvement to one individual’s personal lottery is a pro tanto improvement to the lottery as a whole. But in certain cases, infinitely many pro tanto improvements can add up to an overall worsening—this phenomenon crops up in various well-known infinitary decision problems, like Rouble Trouble and Satan’s Apple (Arntzenius et al. 2004). Table 9 is, we think, another case of this kind. If that’s the right story to tell, then it also casts some doubt on Personal Good in the context of finite-support lotteries over infinite populations—more specifically, for pairs of lotteries such that transforming one into the other would require permutations or other changes to infinitely many personal lotteries. Even if each of these changes in isolation is neutral or positive, their overall effect might be non-positive. Thus, although Negative Dominance and Personal Good are consistent for finite-support lotteries on infinite populations, we think our arguments for completeness are less compelling in that context than in strictly finite contexts.
Objection
Negative Dominance should be rejected, because its appeal is fully captured by the following, weaker principle:
@@
Lottery $L_{1}$ is better than lottery $L_{2}$ only if, on every admissible completion of the personal and impersonal betterness relations, some possible outcome $o_{1}$ of $L_{1}$ is better than some possible outcome $o_{2}$ of $L_{2}$.
An “admissible completion” is one that satisfies some set of fundamental constraints on the ranking of outcomes. Suppose these constraints include Transitivity and Pareto. Then, in the case in Table 3, every admissible completion of the betterness rankings of lives and outcomes must rank some outcome of Lottery 2 (either $(a^{+},b)$ or $(b,a)$) above some outcome of Lottery 1 (either $(a,a)$ or $(b,b)$). [^33] Suppose that, on a given completion, $a$ is at least as personally good as $b$. Then, by Transitivity, $a^{+}$ is better than $b$, so by Pareto, $(a^{+},b)$ is better than $(b,b)$. The other possibility is that $b$ is better than $a$, in which case $(b,a)$ is better than $(a,a)$ by Pareto. Thus, unlike Negative Dominance, Supervaluational Negative Dominance does not rule out that Lottery 2 is better than Lottery 1. More generally, we conjecture that Supervaluational Negative Dominance is consistent, in finite contexts, with the various sets of assumptions that were shown to conflict with Negative Dominance in §§45, in which case a retreat to Supervaluational Negative Dominance would block our arguments for completeness. And one might think that Supervaluational Negative Dominance captures everything that is appealing about Negative Dominance: If every (admissible) way of filling in the gaps in the betterness relations makes some outcome of $L_{1}$ better than some outcome of $L_{2}$, then “one way or another” there will be some outcome-based rationale for the superiority of $L_{1}$, even if the actual, incomplete betterness relation on outcomes does not single out particular outcomes to furnish that rationale.
Response
There are two reasons why we regard Supervaluational Negative Dominance as an inadequate substitute for Negative Dominance. First, it does not capture the motivation for Negative Dominance that we gave in §2. The idea central to our defense of Negative Dominance was that what ultimately matters in evaluating lotteries is how things actually turn out, and how things actually turn out will be a single outcome, not a disjunction or mixture of outcomes. Within our preference-based analysis of the value of lotteries, retreating to Supervaluational Negative Dominance would mean that you can get a requiring reason to prefer $L_{1}$ over $L_{2}$ from the fact that “If you were to fill in the gaps in the betterness relation, then some outcome of $L_{1}$ (either this outcome, or that outcome, or…) would be better than some outcome of $L_{2}$.” It is exactly this appeal to disjunctions of outcomes that, we think, a focus on actual value precludes.
Second, Supervaluational Negative Dominance gives a normative significance to completions that seems unwarranted absent a commitment to completeness. To illustrate what we mean, let’s consider a potential rationale for Supervaluational Negative Dominance from the recent literature. Rabinowicz (2008; 2021) gives a developed theory of value relations, on which value relations in general (on outcomes as well as lotteries) are understood in terms of ideal preference relations. In his 2021, he endorses the following normative constraint on incomplete preferences:
Well-Roundedness
If there is a permissible completion of $P$ [an agent’s preference relation] with respect to $G$ [a set of items in the domain of that preference relation], then $P$ is the intersection of some non-empty set of permissible completions of $P$ with respect to $G$. (Rabinowicz 2021: 219)
This principle appears to support Supervaluational Negative Dominance but not Negative Dominance. To see that it supports Supervaluational Negative Dominance, suppose there is an admissible completion of the personal and impersonal betterness relations on which no outcome of $L_{1}$ is better than any outcome of $L_{2}$. Assuming that an admissible completion of the betterness relation corresponds to a permissible completion of the agent’s preferences, and that all permissible completions of the agent’s preferences will satisfy at least very minimal (“positive”) dominance principles, there will be a permissible completion of the agent’s preferences on which $L_{2}$ is strictly preferred to $L_{1}$, so a preference for $L_{1}$ over $L_{2}$ will not belong to the intersection of the set of permissible completions.
On the other hand, to see that Well-Roundedness does not support Negative Dominance, consider our central example from §§1/4. If the permissible completions of an agent’s preference relation must all satisfy Personal Good (or even a weakening of Personal Good that applies only to personal lotteries over totally ordered lives), then those completions will all include a preference for the sweetened Lottery 2 over the unsweetened Lottery 1. Thus that preference will belong to the intersection of any set of permissible completions. So Rabinowicz’s principle at least leaves open that one is required to prefer Lottery 2. And this would contradict Negative Dominance given Rabinowicz’s (and our) preference-based understanding of the betterness relation on lotteries, together with the auxiliary assumptions of Incomparability Transmission and Small Improvements.
But the principle of Well-Roundedness, we think, gives an unwarranted normative significance to completions. In the cases where it applies (where its antecedent is satisfied), Rabinowicz’s principle can be paraphrased as follows: “Suppose it’s required that, if the agent’s preferences are complete, then they include preference $x$. Then preference $x$ is required simpliciter.” Why accept this principle, if completeness itself is not required? There are many optional features of a preference relation such that any permissible preferences with that feature would include some particular preference. But if the feature itself is optional, then that particular preference may be too.
In short, we find Rabinowicz’s principle of Well-Roundedness less compelling than the premises of our argument for Negative Dominance. More generally, Supervaluational Negative Dominance is too weak to capture the core motivations for Negative Dominance.
@@ 8 Conclusion
Against weak background assumptions, Negative Dominance and Personal Good together rule out incompleteness in the personal value of lives and the impersonal value of finite outcomes and lotteries. We see this as a compelling argument against those forms of incompleteness. Of course, as with any deductive argument, some might find the considerations opposing the conclusion so strong that they find it better to reject an initially plausible premise than to accept the conclusion. But even those who respond in this way to our arguments will have learned something surprising. If they reject Negative Dominance, they will have to give up the very natural idea that the comparative value of lotteries must be explained at least in part by the comparative value of their individual outcomes. And they will have to reject the stronger principle of Statewise Negative Dominance, which has figured centrally in the approaches to risk advocated by many fans of incompleteness (Schoenfield 2014; Bales et al. 2014; Doody 2023). On the other hand, if they reject Personal Good (even when restricted to situations where equality and diversity are not at stake), they will be giving up even the most minimal connection between the overall value of lotteries and their value for particular individuals.
More generally, our arguments show that proponents of standard forms of evaluative incompleteness must choose between the outcome-based perspective on risky choices represented by Negative Dominance, and the “individualist” perspective represented by Personal Good. These perspectives conflict in infinite contexts even without incompleteness. But one might have hoped to confine such conflicts to the paradox-ridden domain of infinity. Incompleteness forces us to abandon one of these perspectives even for strictly finite choices. And, we have argued, this is a substantial cost. [^34] An earlier version of this paper was circulated under the title “Share the Sugar”. For helpful discussion and feedback, we are grateful to Ralf Bader, Pietro Cibinel, Will Combs, Johan Gustafsson, Martin Peterson, Ralph Wedgwood, two anonymous referees, and audiences at the Global Priorities Institute, the 2024 meeting of the Philosophy, Politics, and Economics Society, Texas A&M, the University of Toronto, and the University of Southern California.